OpenStudy (anonymous):

evaluate: (1/3b)^-2 for b =12

OpenStudy (callisto):

\[(1/3b)^{-2} \]Put b=12 \[(1/[3(12)])^{-2} = (1/36)^{-2} = \frac{1}{(\frac{1}{36})^2} = 36^2 =?\]

OpenStudy (anonymous):

1269